OpenStudy (anonymous):

What is the volume of a sphere with a surface area of 9pi yd^2?

OpenStudy (anonymous):

surface area of a sphere: A=4pi*r^2 Volume: V=4/3 pi*r^3 you know surface area - using to substitute known values in V

OpenStudy (anonymous):

Surface Area of a Sphere = 4*pi*r^2 so 4*pi*r^2 = 9*pi pi cancels r^2 = 9/4 r = \[\sqrt{9/4}\] r = 1.5 yd